# CLOT: Closed Loop Optimal Transport for Unsupervised Action Segmentation

### v1

@@ Abstract
Unsupervised action segmentation has recently pushed its limits with ASOT, an optimal transport (OT)-based method that simultaneously learns action representations and performs clustering using pseudo-labels. Unlike other OT-based approaches, ASOT makes no assumptions on the action ordering, and it is able to decode a temporally consistent segmentation from a noisy cost matrix between video frames and action labels. However, the resulting segmentation lacks segment-level supervision, which limits the effectiveness of the feedback between frames and action representations. To address this limitation, we propose Closed Loop Optimal Transport (CLOT), a novel OT-based framework that introduces a multi-level cyclic feature learning mechanism. Leveraging its encoder-decoder architecture, CLOT learns pseudo-labels alongside frame and segment embeddings by solving two separate OT problems. It then refines both frame embeddings and pseudo-labels through cross-attention between the learned frame and segment embeddings, integrating a third OT problem. Experimental results on four benchmark datasets demonstrate the benefits of cyclical learning for unsupervised action segmentation.  [^1] https://github.com/elenabbbuenob/CLOT
@@ 1 Introduction
Unsupervised action segmentation, the task of labeling each frame in a video or a set of videos with its action class without relying on a labeled training set, has been gaining increasing attention in recent years [12, 23, 28, 41, 51, 45]. This task is critical for numerous real-world applications, including action detection in sports events, video surveillance, and robotic manipulation [18, 54]. The classical approach to unsupervised action segmentation follows a multi-step pipeline, where frame embeddings are first learned and then clustered into action segments [22, 47, 26]. More recently, optimal transport (OT)-based methods have emerged, jointly learning action representations and clustering frames by leveraging OT theory to generate pseudo-labels for self-training [51, 23, 45].
@@
Figure 1: Given precomputed frame embeddings $X$ as input, CLOT first solves an OT problem to generate pseudo-labels $T$ alongside initial frame embeddings $F$. It then estimates segment embeddings $S$ and pseudo-labels $T_{S}$ through a second OT problem. Finally, a third OT problem is employed to obtain refined frame embeddings $F_{R}$ and pseudo-labels $T_{R}$, leveraging cross-attention between the learned $F$ and $S$, thereby closing the loop between frame and segment representations.
@@
Unlike previous methods, our approach, illustrated in Fig. 1, builds upon the strengths of OT-based methods but extends prior formulations by incorporating a multi-level cyclic feature learning mechanism. Initially, it generates pseudo-labels and jointly learns frame embeddings and segment embeddings by solving two separate OT problems. To further improve segmentation consistency, CLOT refines the pseudo-labels through a cross-attention mechanism between frame and segment representations, integrating a third OT problem into the process.
Our cyclic optimization framework runs on top of a carefully designed encoder-decoder-based architecture (see Fig. 2), consisting of an MLP-based encoder endowed with a feature dispatching mechanism, a parallel decoder and a cross-attention-based mechanism. In addition, our approach leverages the Sliced Wasserstein (SW) distance to refine the initial cost matrix derived from frame embeddings. Our contributions are as follows:
@@
    We propose a novel framework for unsupervised action segmentation that enforces segment-level consistency, achieving fine-grained segmentation boundaries while ensuring generalization across different videos.
@@
    We achieve very competitive performance at both video-level and activity-level over four benchmark datasets, namely Breakfast [21], YouTube INRIA [2], 50salads [42] and Desktop Assembly [23].
@@ 2 Related work
Recently, Optimal Transport (OT) has emerged as a powerful tool for jointly learning action representations and pseudo-labels through self-training, enabling effective feedback between representation learning and clustering while directly optimizing for action segmentation. The pioneering work TOT [23] introduced a temporal OT formulation that generates pseudo-labels from predicted cluster assignments. However, this approach assumes a fixed action ordering across all videos within an activity and uniform label assignment, contradicting the natural variability and long-tailed distribution of action labels. UFSA [45] addressed these limitations by incorporating frame- and segment-level cues from transcripts, allowing for action permutations within activities and non-uniform label assignments. However, it still requires prior knowledge of an estimated action order to infer the segmentation.
@@
Figure 2: Diagram of the CLOT architecture. The initial feature embeddings, $X$, are extracted from video frames, while $A$ represents the action embeddings. The model consists of a multi-layer perceptron (MLP) encoder with a feature dispatching mechanism, a parallel decoder, and a cross-attention module. The purple boxes indicate the architectural components of the model, whereas the orange boxes correspond to the learnable variables. Arrows denote the computational/gradient flow, with crossed-out arrows ($\rightarrow$) indicating points where the gradient flow is stopped. Specifically, black arrows correspond to the encoder, green arrows to the decoder, and purple arrows to the cross-attention stage. The model leverages a closed-loop optimal transport mechanism to refine frame embeddings through cross-attention with segment embeddings, ensuring improved action segmentation accuracy. The estimated ($\mathbf{T},\mathbf{T}_{R},\mathbf{T}_{S}$) are used as a pseudo-label to compute the losses in each stage.
@@ 3.2 Optimal Transport formulation
[…]
@@ 4 Experimental Results
Table 2: Comparisons on the Breakfast [21], Youtbe Instr. [2], 50Salads [42] and Desktop Assembly [23] datasets computed by applying Hungarian matching per video basis. * denotes whether the method has a training stage on target videos. The highest accuracy is indicated in bold, and the second highest accuracy is underlined.
@@ 4.2 State-of-the-Art Comparisons
Traditionally, unsupervised action segmentation at the video level and the activity level have been treated as separate challenges. Video-level segmentation is generally considered the less complex of the two, as it does not require generalizing action representations across multiple videos. In contrast, activity-level segmentation demands the ability to maintain accurate action boundary detection while ensuring generalization across different instances of the same activity. This trade-off has historically limited performance improvements across both settings. CLOT establishes a new state-of-the-art in both video-level and activity-level segmentation. It consistently outperforms previous approaches on almost all benchmark datasets. Below, we analyze our results in detail, comparing them to existing methods and interpreting the architectural contributions that drive our superior performance. Fig 4 contains qualitative examples where actions are seen in different orders across videos within a dataset. More qualitative results are in the supp. mat.
Figure 4: Qualitative results. We display the ground-truth (GT), the results of CLOT (Ours) and ASOT [51]. Results from different datasets, activities and observation horizons are displayed for comparison.
@@
CLOT surpasses all existing methods, achieving a significant performance gain. Specifically, our approach improves over the previous best-performing method by an average of $+2.54\%$ in MoF, $+5.08\%$ in F1 Score, and $+3.2\%$ in mIoU across five benchmark datasets.
@@
CLOT outperforms ASOT [51] across all datasets except for YTI, while often achieving F1 and mIoU scores comparable to methods specifically designed for video-level segmentation. This improvement suggests that CLOT dynamically refines action representations, allowing it to adapt more effectively to varying segment structures, which is particularly important at video-level.
@@ Acknowledgements
This work was supported by grant PRE2020-094714 and project PID2023-151351NB-I00 funded by MCIN/ AEI /10.13039/501100011033, by ”ESF Investing in your future” and by ERDF, UE.
Supplementary Material
@@
Institut de Robòtica i Informàtica Industrial, CSIC-UPC
Barcelona, Spain
@@
This supplementary material provides extended analyses and implementation details to support our proposed CLOT paper. It includes sensitivity studies on key parameters, complexity comparisons with ASOT, and additional ablations to isolate the effect of each component. We also provide full hyperparameter settings and extra qualitative results across all datasets to further illustrate the advantages of our approach.
@@ Appendix A Sensitivity Analysis
[…]
@@ Appendix B Complexity
CLOT introduces a multi-level cyclic learning mechanism, solving three OT problems instead of one, which could lead to increased computational costs. However, to maintain efficiency, we incorporate an entropy-regularized solver, similar to ASOT [51], which accelerates the OT optimization process while maintaining accuracy. This approach leverages sparsity structures in the cost matrices, enabling faster convergence with a computational complexity of $O(NK)$ per iteration. As a result, CLOT remains scalable to large datasets, efficiently handling videos with thousands of frames while keeping processing times competitive. The overall runtime remains comparable to ASOT, with only a moderate increase in memory usage due to additional attention layers in the decoder. Tab. 4 shows the GFLOPs and per-batch training time, using flop counter in PyTorch. Despite CLOT using 3 OT, and adding complexity to the architecture, it increases FLOPs and runtime moderately.
@@ Appendix C Additional Ablation Study
To assess the influence of decoding strategies on segmentation quality, we compare our parallel decoder with its autoregressive counterpart similar to [46] across both activity-level and video-level evaluations (Tab 5). The parallel design, aligned with CLOT’s architecture, avoids error propagation by predicting segments simultaneously rather than sequentially. Results indicate consistent improvements across all datasets and metrics, with notable gains in F1 and mIoU, especially for short or overlapping actions, highlighting the benefit of structured segment-level modelling over temporally fragile autoregressive predictions.
Comparison to ASOT.
ASOT corresponds to our 1st stage (frame-to-action) without SWD and FD.These are not plug-and-play modules but part of a coherent architectural redesign. In Fig. 6, both components bring measurable benefits over ASOT, while their absence in CLOT underscores the critical role they play in overall performance.
@@ Appendix D Implementation Details
Appendix E Additional Qualitative Results
Figure 7: Qualitative results. We display the ground-truth (GT), the results of CLOT (Ours) and ASOT [51]. Results from different datasets, activities and observation horizons are displayed for comparison.
For a more comprehensive analysis, we present additional qualitative results from all four datasets in Fig. 7. These visualizations illustrate how CLOT effectively refines segmentation boundaries compared to ASOT and other baseline methods.

### v2

@@ Abstract
Unsupervised action segmentation has recently pushed its limits with ASOT, an optimal transport (OT)-based method that simultaneously learns action representations and performs clustering using pseudo-labels. Unlike other OT-based approaches, ASOT makes no assumptions about action ordering and can decode a temporally consistent segmentation from a noisy cost matrix between video frames and action labels. However, the resulting segmentation lacks segment-level supervision, limiting the effectiveness of feedback between frames and action representations. To address this limitation, we propose Closed Loop Optimal Transport (CLOT), a novel OT-based framework with a multi-level cyclic feature learning mechanism. Leveraging its encoder-decoder architecture, CLOT learns pseudo-labels alongside frame and segment embeddings by solving two separate OT problems. It then refines both frame embeddings and pseudo-labels through cross-attention between the learned frame and segment embeddings, by integrating a third OT problem. Experimental results on four benchmark datasets demonstrate the benefits of cyclical learning for unsupervised action segmentation.  [^1] https://github.com/elenabbbuenob/CLOT
@@ 1 Introduction
Unsupervised action segmentation, the task of labeling each frame in a video or a set of videos with its action class without relying on a labeled training set, has been gaining increasing attention in recent years [12, 23, 28, 41, 51, 45]. This task is critical for numerous real-world applications, including action detection in sports events, video surveillance, and robotic manipulation [18, 54]. The classical approach to unsupervised action segmentation follows a multi-step pipeline, where frame embeddings are first learned and then clustered into action segments [22, 47, 26]. More recently, optimal transport (OT)-based methods have emerged, jointly learning action representations and clustering frames by leveraging OT theory to generate pseudo-labels for self-supervised training. [51, 23, 45].
@@
Figure 1: Given precomputed frame embeddings $X$ as input, CLOT first solves an OT problem to generate pseudo-labels $T$ alongside initial frame embeddings $F$. It then estimates segment embeddings $S$ and pseudo-labels $T_{S}$ by solving a second OT problem. Finally, a third OT problem is formulated to obtain refined frame embeddings $F_{R}$ and pseudo-labels $T_{R}$, by leveraging cross-attention between the learned $F$ and $S$, thereby closing the loop between frame and segment representations.
@@
Our approach, as illustrated in Fig. 1, initially generates pseudo-labels and jointly learns the frame embeddings. These embeddings are then used to produce segment embeddings. Both are optimized by solving two separate OT problems. To further improve segmentation consistency, CLOT refines the pseudo-labels through cross-attention between frame and segment representations, by integrating a third OT problem.
Our cyclic optimization framework employs a carefully designed encoder-decoder architecture (see Fig. 2), comprising an MLP-based encoder equipped with a feature dispatching mechanism, a parallel decoder, and a cross-attention-based mechanism. Additionally, we incorporate the Sliced Wasserstein (SW) distance to refine the initial cost matrix, leveraging its efficiency and robustness in high-dimensional representation spaces. Our contributions are as follows:
@@
    We propose a novel framework for unsupervised action segmentation that enforces segment-level consistency, achieving fine-grained segmentation and generalization across different videos.
@@
    We achieve very competitive performance at both video-level and activity-level over four benchmark datasets, namely Breakfast [21], YouTube INRIA [2], 50salads [42], and Desktop Assembly [23].
@@ 2 Related work
Recently, Optimal Transport (OT) has emerged as a powerful tool for jointly learning action representations and pseudo-labels through self-training, enabling effective feedback between representation learning and clustering while directly optimizing for action segmentation. TOT [23] introduced a temporal OT formulation that generates pseudo-labels from predicted cluster assignments. However, this approach assumes fixed action ordering across all videos and uniform label assignment, contradicting the natural variability and long-tailed distribution of action labels. UFSA [45] addressed these limitations by incorporating frame- and segment-level cues from transcripts, allowing for action permutations within activities and non-uniform label assignments. However, it still requires prior knowledge of an estimated action order to infer the segmentation.
@@
Figure 2: Diagram of the CLOT architecture. The only model input is the initial feature embeddings, $X$, extracted from video frames. The model consists of a multi-layer perceptron (MLP) encoder with a feature dispatching mechanism, a parallel decoder, and a cross-attention module. The magenta boxes indicate the architectural components of the model, whereas the orange boxes correspond to the learnable variables, which include the learnable action embeddings AA as well as the learnable frame and segment embeddings. Arrows denote the computational/gradient flow, with crossed-out arrows ($\rightarrow$) indicating points where the gradient flow is stopped. Specifically, cyan arrows correspond to the first stage, green arrows to the second stage, and purple arrows to the third stage. The model leverages a closed-loop optimal transport mechanism to refine frame embeddings through cross-attention with segment embeddings, ensuring improved action segmentation accuracy. The estimated OT matrices ($\mathbf{T},\mathbf{T}_{R},\mathbf{T}_{S}$) act as pseudo-labels during training, and are computed at different stages by using the frame/segment features ($F$, $S$, and $F_{R}$) together with the action embeddings $A$, which are used to define the OT cost matrices.
@@ 3.2 Optimal Transport formulation
$W_{p}(\mu,\nu)=\left(\inf_{\mathcal{T}\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}d^{p}(x,y)d\mathcal{T}(x,y)\right)^{\frac{1}{p}},$ (8)
@@ 4 Experimental Results
Table 2: Comparisons on the Breakfast [21], Youtbe Instr. [2], 50Salads [42] and Desktop Assembly [23] datasets computed by applying Hungarian matching per video. * denotes whether the method has a training stage on target videos. The highest accuracy is indicated in bold, and the second highest is underlined.
@@ 4.2 State-of-the-Art Comparisons
Traditionally, unsupervised action segmentation at the video level and the activity level have been treated as separate challenges. Video-level segmentation is generally considered the less complex of the two, as it does not require generalizing action representations across multiple videos. In contrast, activity-level segmentation demands the ability to maintain accurate action boundary detection while ensuring generalization across different instances of the same activity. This trade-off has historically limited performance improvements across both settings. CLOT establishes a new state-of-the-art in both video-level and activity-level segmentation. It consistently outperforms previous approaches on almost all benchmark datasets. Below, we analyze our results in detail, comparing them to existing methods and interpreting the architectural contributions that drive our superior performance. Fig 4 contains qualitative examples where actions are seen in different orders across videos within a dataset. More results are in the supp. mat.
Figure 4: Qualitative results. We display the ground-truth (GT), the results of CLOT (Ours), and ASOT [51]. Results from different datasets and activities are displayed for comparison.
@@
CLOT surpasses all existing methods, achieving a significant performance gain. Specifically, our approach improves over the previous best-performing method by an average of $2.54\%$ in MoF, $5.08\%$ in F1 Score, and $3.2\%$ in mIoU across five benchmark datasets.
@@
CLOT and ASOT [51] learns a single model across all videos of an activity. In contrast, * methods are trained and optimized independently per video and are tailored to each test instance, often achieving higher performance by design, without generalization across videos. CLOT outperforms ASOT [51] across all datasets except for YTI, while often achieving F1 and mIoU scores comparable to methods specifically designed for video-level segmentation. This improvement suggests that CLOT dynamically refines action representations, allowing it to adapt more effectively to varying segment structures, which is particularly important at the video-level.
@@ Acknowledgements
This work was supported by grant PRE2020-094714, the project PID2019-110977GA-I00 and the project PID2023-151351NB-I00, funded by Ministerio de Ciencia e Innovación (MCIN)/ Agencia Estatal de Investigación (AEI) /10.13039/501100011033, by European Social Fund (ESF) Investing in your future and by ERDF, UE. It was also supported by an Alexander von Humboldt (AvH) fellowship for experienced researchers funded by the AvH Foundation.
Supplementary Material for
CLOT: Closed Loop Optimal Transport for Unsupervised Action Segmentation
@@
Institut de Robòtica i Informàtica Industrial, CSIC-UPC, Barcelona, Spain
@@
This supplementary material provides extended analyses and implementation details to support our proposed CLOT paper. It includes sensitivity studies on key parameters, complexity comparisons with ASOT, and additional ablations to isolate the effect of each component. We also provide full hyperparameter settings and additional qualitative results across all datasets to illustrate the advantages of our approach further.
@@ Appendix A Sensitivity Analysis
$K^{\prime}$ controls the capacity of the decoder, that is the maximum number of predicted segments, the actual number is learned. We set $K^{\prime}=K+\text{nseg}$, where $K$ is the number of action classes and $\text{nseg}\in\mathbb{Z}$ is a dataset-specific offset. See the results for $\text{nseg}\in[-6,+6]$ in Fig. 5 (first three tables). The straight lines indicate the SOTA. CLOT achieves equal or superior performance in F1 and mIoU across nearly all nseg in BF, 50S(Mid), 50S(Eval), and DA. For MoF, CLOT consistently outperforms or equals the SOTA $\forall\text{nseg}$ in BF, 50S (Mid), and YTI. In particular, negative nseg can help reduce over-segmentation in presence of background or very unbalanced action durations.
@@
Figure 5: Sensitivity analysis for MoF, IoU and F1 on nseg and and MoF on $P$.
DA ($\circ$, solid), FS (Mid) ($\square$, dashed), FS (Eval) ($\triangle$, dotted), BF ($\Diamond$, dash-dot), and YTI ($\triangledown$, custom dash).
@@ Appendix B Complexity
CLOT introduces a multi-level cyclic learning mechanism, solving three OT problems instead of one, which could lead to increased computational costs. However, to maintain efficiency, we incorporate an entropy-regularized solver, similar to ASOT [51], which accelerates the OT optimization process while maintaining accuracy. This approach leverages sparsity structures in the cost matrices, enabling faster convergence with a computational complexity of $O(NK)$ per iteration. As a result, CLOT remains scalable to large datasets, efficiently handling videos with thousands of frames while keeping processing times competitive. The overall runtime remains comparable to ASOT, with only a moderate increase in memory usage due to additional attention layers in the decoder. Tab. 4 shows the GFLOPs and per-batch training time, using flop counter in PyTorch. Despite CLOT using 3 Optimal Transport (OT) problems and adding architectural complexity, it increases Floating Point Operations (FLOPs) and runtime moderately.
@@ Appendix C Additional Ablation Study
To assess the influence of decoding strategies on segmentation quality, we compare our parallel decoder with its autoregressive counterpart, similar to [46] across both activity-level and video-level evaluations (Tab 5). The parallel design, aligned with CLOT’s architecture, avoids error propagation by predicting segments simultaneously rather than sequentially. Results indicate consistent improvements across all datasets and metrics, with notable gains in F1 and mIoU, especially for short or overlapping actions, highlighting the benefit of structured segment-level modelling over temporally fragile autoregressive predictions.
@@
Comparison to ASOT.
ASOT corresponds to our 1st stage (frame-to-action) without SWD and FD. However, it is not meaningful to treat SWD or FD as independent add-ons to ASOT, as our method redefines the OT formulation itself by integrating both components to enhance representations prior to transport resolution. These are not plug-and-play modules but part of a coherent architectural redesign. In Fig. 6, both components bring measurable benefits over ASOT, while their absence in CLOT underscores the critical role they play in overall performance.
@@ Appendix D Implementation Details
Appendix E Additional Results
Quantitative Results per Activity
To provide a more comprehensive evaluation, we report class-wise performance for the Breakfast and YTI datasets in Tab. 6 across key metrics such as MOF, F1, and mIoU.
Qualitative Results
Figure 7: Qualitative results. We display the ground-truth (GT), the results of CLOT (Ours), and ASOT [51]. Results from different datasets and activities for comparison.
We also present additional qualitative examples across all four datasets in Fig. 7. These visualizations demonstrate CLOT’s ability to more accurately delineate action boundaries when compared to ASOT and other baseline methods, highlighting its superior segmentation performance.
Breakfast Dataset
Activity MoF F1 mIoU
coffee 39.4 26.3 17.0
cereals 60.4 51.1 29.0
tea 58.2 48.0 24.0
milk 51.0 48.0 24.9
juice 73.5 46.9 56.3
sandwich 67.9 37.6 53.4
scrambledegg 51.2 34.1 52.2
friedegg 65.1 39.1 48.2
salat 43.7 41.2 57.2
pancake 68.3 29.5 36.3
YTI Dataset
Activity MoF F1 mIoU
changing tire 49.9 30.3 21.0
coffee 67.2 44.3 30.9
cpr 60.2 46.8 23.0
jump car 35.7 26.3 17.0
repot 43.7 27.2 14.0
Table 6: Per-class performance on the Breakfast [21] and Youtube Instr.  [2] datasets.